OpenStudy (anonymous):

tan θ/(sec θ − cos θ) Write the trigonometric expression in terms of sine and cosine, and then simplify.

OpenStudy (anonymous):

(sin θ/cos θ)/(1/cos θ)-cos θ is what I have in terms of sin and cosine, I'm not really sure how to simplify this.

OpenStudy (anonymous):

(sinθ/cosθ)cosθ-cosθ =sinθ-cosθ

OpenStudy (anonymous):

The final answer should be csc θ. There is a division sign inbetween the (sin θ/cos θ) and (1/cos θ) - cos θ.

OpenStudy (anonymous):

ahh, so it's: tan θ/(sec θ − cos θ) ?

OpenStudy (ajprincess):

Is your question this tanx/(secx-cosx)

OpenStudy (anonymous):

yea sorry i didnt put the parenthesis in the original question.

OpenStudy (anonymous):

sec θ − cos θ = 1/cos θ -cos θ =(1- cos^2 θ)/cos θ = sin^2θ/cos θ Now: tanθ /sin^2θ/cos θ=( sinθ/cosθ)(cosθ /sin^2θ) =1/sinθ= coscθ

OpenStudy (ajprincess):

sinx/cosx((1/cosx)-cosx) =sinx/cosx((1-cos^2x)/cosx) =sinx/sin^2x =1/sinx =cosecx 1-cos^2x=sin^2x This is derived from sin^2x+cos^2x=1

OpenStudy (anonymous):

\[(\sin \theta /\cos \theta)/((1/\cos \theta)-\cos \theta)\] on simplification \[\sin \theta/1-\cos^2\theta\] \[\sin \theta/\sin^2\theta\] \[1/\sin \theta\] \[cosec \theta\]